### v1

# Single impurity atom embedded in a dipolar two-soliton molecule as a qubit

Abstract
We consider a single impurity atom trapped in a double well (DW) potential created by a dipolar two-soliton molecule in a quasi-one-dimensional geometry. By solving the eigenvalue problem for the impurity atom in the DW potential, we find that its ground and first excited states are well separated from higher excited states. This allows it to be approximated by a desirable two-level quantum system. Numerical simulations of the Schrödinger equation, governing impurity atom, demonstrate periodic oscillations in the probability of finding the impurity confined either to the “left” or to the “right” side of the DW potential. An analytic expression for the coherent oscillations of the population imbalance between the two wells of the DW potential has been derived using the two-mode approximation. Theoretical predictions of the mathematical model are in good agreement with the results of numerical simulations. Potential usage of the developed setup as a physical realization of “qubit” has been discussed.
I Introduction
The enhanced computational power of quantum computers will be primarily due to the use of quantum bits called qubit. Unlike its classical counterpart, called bit, which enables operations using only two digits “0” and “1”, qubits can exist in multiple states simultaneously, owing to the superposition principle of quantum mechanics, thus leading to an enormous increase in speed of operations. To date, the qubits have been realized on different physical platforms, such as Josephson junctions, quantum dots, nuclear magnetic resonances, trapped ions, neutral atoms in optical tweezers, cold atoms in optical lattice, coherent photons, etc. (see e.g. book nakahara-book ). Each of these realizations has distinct advantages and limitations, and lack of a universally accepted platform motivates the search for alternative realizations. Impurity atoms confined to matter-wave bright and dark solitons in the context of their qubit properties were investigated in Refs. shaukat2017 ; shaukat2019 . The key benefit of this approach is that the qubit is well protected from decoherence salerno2005 . Previous research primarily examined single hump matter-wave solitons with an embedded impurity atom. The energy levels of a particle in a single well potential, such as Gaussian nandi2010 or Pöschl-Teller type flugge-book ; senn1986 , are well known to be considerably spaced apart, making them less suitable for the functioning of a qubit. In contrast, double-well potentials exhibit closely spaced energy levels, making them suitable for qubit representation foot2011 . In this context, tunnel-coupled semiconductor double quantum dots were considered highly promising systems divincenzo2005 ; wiel2003 .
Here, we propose a qubit realization using a single impurity atom embedded in a dipolar two-soliton molecule that forms a self-induced double well (DW) potential. This configuration enables a robust two-level qubit system, while the dipolar nonlocality offers in-situ tunability of the barrier height/width and energy splitting through the strength of inter-atomic interactions. It is important to note that matter wave solitons in Bose-Einstein condensates (BEC) with only contact atomic interactions do not support stable bound states resembling molecules. Therefore, ordinary BECs cannot provide a stable DW potential for the impurity atom.
Our strategy in this work goes as follows: we first develop a variational approach to find the stationary state of the dipolar two-soliton molecule, then numerically calculate the energy spectrum of the impurity atom trapped in the molecular potential. In the framework of the two-mode approximation we derive the frequency of coherent oscillations of the population imbalance between the two wells of the DW potential. To validate our findings, the analytical predictions will be compared with numerical simulations. Finally, we discuss the potential usage of our results in the design a soliton-based qubit.
II The wave profile of the two-soliton molecule
The model is described by a 1D nonlocal GPE supporting a two-soliton molecule and a linear Schrödinger equation for a single impurity atom embedded in the molecular potential
$i\psi_{t}$ $=$ $-\frac{1}{2}\psi_{xx}-q|\psi|^{2}\psi-g\psi\int^{\infty}_{-\infty}R(|x-x^{\prime}|)|\psi(x^{\prime},t)|^{2}dx^{\prime},$ (1)
$i\phi_{t}$ $=$ $-\frac{1}{2}\phi_{xx}-\gamma|\psi|^{2}\phi,$ (2)
where $\psi(x,t)$ is the mean field wave function of the dipolar soliton molecule, $\phi(x,t)$ is the wave function of the impurity atom, $q$ and $g$ are the strengths of contact and dipole-dipole attractive forces between atoms of the condensate. In experiments, both of these parameters can be tuned, the former by a Feschbach resonance method chin2010 , and the latter using rotating magnetic fields tang2018 . In this setting, the impurity “feels” the presence of a double well potential due to the dipolar two-soliton molecule with a coupling strength $\gamma$.
The long-range interactions between atoms of the dipolar BEC is characterized by a normalized kernel function, chosen to be of Gaussian type
$R(x)=\frac{1}{\sqrt{2\pi}w}\exp\left(-\frac{x^{2}}{2w^{2}}\right),\qquad\int_{-\infty}^{+\infty}R(x)dx=1,$ (3)
where $w$ is a parameter that characterizes the nonlocal nature of the dipolar atomic interactions in BEC. This expression is selected for a computational convenience, although it is more relevant to optical media. Qualitative similarity between the Gaussian kernel function (3) and other expressions used in BEC research, such as the single mode approximation kernel sinha2007 and that with a cutoff parameter cuevas2009 , justifies our choice.
The objective of this work is to develop an analytical framework for the “qubit problem”, based on the variational approach (VA). The validity of the VA for describing the properties of dipolar BEC over a wide range of experimental parameters was demonstrated in yi2001 ; olson2013 . It is evident from the basic model Eqs. (1)-(2) that the GPE is decoupled from the Schrödinger equation, while the impurity “feels” the dipolar BEC only as an external potential. Obviously, a single impurity atom cannot significantly distort the wave profile of the BEC, which consists of a large number of dipolar atoms. Similar models have been employed, and proven to be realistic, in previous publications devoted to BEC-based qubits shaukat2017 ; shaukat2019 ; salerno2005 .
The wave profile of the two-soliton molecule in a dipolar BEC, that is, the double well potential acting on the impurity, can be determined using VA. To this end, we write the Lagrangian density corresponding to GPE (1)
${\cal L}=\frac{i}{2}(\psi\psi^{\ast}_{t}-\psi^{\ast}\psi_{t})+\frac{1}{2}|\psi_{x}|^{2}-\frac{q}{2}|\psi|^{4}-\frac{g}{2}|\psi|^{2}\int\limits_{-\infty}^{\infty}R(x-x^{\prime})|\psi(x^{\prime},t)|^{2}dx^{\prime},$ (4)
A suitable trial function for the dipolar two-soliton molecule was shown to have the form baizakov2015
$\psi(x,t)=A(t)\,x\,\exp\left[-\frac{x^{2}}{2a(t)^{2}}+ib(t)x^{2}+i\varphi(t)\right],$ (5)
where $A(t)$, $a(t)$, $b(t)$ and $\varphi(t)$ are variational parameters, associated with the amplitude, width, chirp and phase, respectively. The norm $N=\int|\psi(x)|^{2}dx=A^{2}a^{3}\sqrt{\pi}/2$ is proportional to the number of atoms in the dipolar condensate and it represents the conserved quantity of GPE, (1). Substitution of the ansatz (5) and response function (3) into the Lagrangian density (4) and subsequent integration over the space variable $x$ yields the effective Lagrangian
$\frac{L}{N}=\frac{3}{4a^{2}}+3a^{2}b^{2}+\frac{3}{2}a^{2}b_{t}+\varphi_{t}-\frac{3qN}{8\sqrt{2\pi}a}-\frac{gN}{8\sqrt{2\pi}}\frac{3a^{4}+4a^{2}w^{2}+4w^{4}}{(a^{2}+w^{2})^{5/2}}.$ (6)
The VA equation for the parameter $a$ of the two-soliton molecule can be derived from the corresponding Euler-Lagrange equations $d/dt(\partial L/\partial\nu_{t})-\partial L/\partial\nu=0$ for variational parameters $\nu\rightarrow a,b,\varphi$, using the effective Lagrangian (6)
$a_{tt}=\frac{1}{a^{3}}-\frac{qN}{4\sqrt{2\pi}a^{2}}-\frac{gN}{4\sqrt{2\pi}}\frac{a(a^{4}+4w^{4})}{(a^{2}+w^{2})^{7/2}}.$ (7)
This equation is similar to the equation of motion of a unit mass particle in the anharmonic potential
$a_{tt}=-\partial U/\partial a\quad\mbox{with}\quad U(a)=\frac{1}{2a^{2}}-\frac{qN}{4\sqrt{2\pi}a}-\frac{gN}{12\sqrt{2\pi}}\frac{3a^{4}+4a^{2}w^{2}+4w^{4}}{(a^{2}+w^{2})^{5/2}},$ (8)
which is plotted in Fig. 1a for particular parameter settings. The minimum of the potential corresponds to the stationary width of the soliton molecule $a_{0}$, whose value can be found from the fixed point of Eq. (7), defined by $a_{tt}=0$, for particular parameter values $g$, NN and $w$. The separation between center-of-mass positions of two anti-phase bright solitons constituting the molecule is $\xi=4a_{0}/\sqrt{\pi}$. At larger separation ($a>a_{0}$) the solitons attract each other ($\partial U/\partial a>0$), and at smaller separation ($a<a_{0}$) they repel ($\partial U/\partial a<0$), therefore the effective potential $U(a)$ has a binding nature necessary for the formation of the stable molecule baizakov2019 . The wave profile of the dipolar two-soliton molecule, numerically constructed using the imaginary time propagation method chiofalo2000 applied to GPE (1), and as predicted by VA, is shown in Fig. 1b.
$\qquad\qquad(a)$                       $(b)$
Refer to caption Refer to caption
Figure 1: (a) The effective potential given by Eq. (8). (b) The wave profile of the two-soliton molecule obtained by numerical solution of the GPE (red line) and according to VA (blue dashed line). Parameter values $A=0.177$, $a=3.299$ follow from VA Eq. (7) for $q=1$, $g=6$, $w=5$, $N=1$.
III Bound states of the impurity atom trapped by a two-soliton molecule
Once the wave profile of the two-soliton molecule has been determined, the bound states of the impurity atom subject to the molecular potential can be identified from the eigenvalue problem associated with Eq. (2)
$H\phi_{j}(x)=\mu_{j}\phi_{j},\qquad j=0,1,2,$..., (9)
where $H=-\frac{1}{2}\partial_{x}^{2}+V(x)$ is the Hamiltonian, and $V(x)=-\gamma|\psi(x)|^{2}$ represents the double-well potential created by the dipolar two-soliton molecule. For numerical implementation, we discretize the spatial domain $x\in[-L/2,L/2]$ and utilize the matrix representation of Eq. (1). Then, using the standard procedure Eigensystem[H] of the Mathematica package, we obtain bound state energies and corresponding wave functions of the impurity atom in the DW potential.
The ground state ($\mu_{0}=-0.326$) and first excited state ($\mu_{1}=-0.312$) wave profiles of the system (9) are shown in Fig. 2a. Evidently, the energy levels of the symmetric ground state $\phi_{0}(-x)=\phi_{0}(x)$, and antisymmetric first excited state $\phi_{1}(-x)=-\phi_{1}(x)$ are close. In addition the doublet components ($\phi_{0},\phi_{1}$) are real and orthonormal with energy splitting $\Delta=\mu_{1}-\mu_{0}>0$
$\int_{-\infty}^{\infty}\phi_{j}(x)\phi_{k}(x)dx=\delta_{j,k},\quad j,k=0,1.$ (10)
The next excited states with a negative energy $\mu_{2}=-0.060$ and $\mu_{3}=-0.018$ are well separated from the lowest two states. This indicates that the dipolar two-soliton molecule with an embedded impurity atom can be approximated as a two-level system. Since any two-level quantum system can function as a qubit, thus the proposed model is appropriate for our purpose.
$\qquad\qquad(a)$                       $(b)$
Refer to caption Refer to caption
Figure 2: (a) The wave profiles of the symmetric ground state with $\mu_{0}=-0.326$ (blue), and the antisymmetric first excited state with $\mu_{1}=-0.312$ (yellow) of the impurity atom in the double well potential. (b) Localized combinations constructed from the linear superposition of the ground and first excited states of the system according to Eq. (11). For visual convenience the double well potential $V(x)=-\gamma|\psi(x)|^{2}$ with $\gamma=4$ is also shown by a dashed black line. The parameter values are similar to Fig. 1.
To observe coherent Josephson type oscillations of the impurity atom in the molecular double well potential we construct the localized combinations (see Fig. 2b)
$\phi_{R}(x)=\frac{\phi_{0}(x)+\phi_{1}(x)}{\sqrt{2}},\qquad\phi_{L}(x)=\frac{\phi_{0}(x)-\phi_{1}(x)}{\sqrt{2}},$ (11)
whose orthonormality can easily be verified. Physically, $\phi_{R}$ ($\phi_{L}$) are confined in the right (left) well of the DW potential because the symmetric ($\phi_{0}$) and antisymmetric ($\phi_{1}$) waves localize when combined as indicated in Eq. (11).
When the initial state for Eq. (2) is selected as corresponding to $\phi_{R}$ or $\phi_{L}$ in Eq. (11), the two-level system evolves smoothly in time such that the probability of finding the atom in the left or right well of the DW potential periodically changes. The frequency of these oscillations is known as a Rabi frequency griffiths-book . External driving fields at Rabi frequency are used for calibration and manipulation of qubits. It is essential to note that the system always remains in a superposition of the two basis states $\phi_{0},\phi_{1}$}, whose probability amplitudes change continuously unless the decoherence mechanisms, including the attempt to measure the system’s state, come into play.
IV Dynamics of the impurity atom in a double-well potential
The time evolution of the system can be described by the celebrated two-mode model smerzi1997 . To this end, we approximate the full wavefunction by projecting it onto the subspace spanned by the lowest energy modes $\{\phi_{L},\phi_{R}\}$
$\phi(x,t)\approx c_{L}(t)\,\phi_{L}(x)+c_{R}(t)\,\phi_{R}(x),$ (12)
where $c_{L}(t),c_{R}(t)$ are complex time-dependent amplitudes. The normalization condition is given by
$\int|\phi(x,t)|^{2}dx\approx|c_{L}|^{2}+|c_{R}|^{2}=1.$ (13)
In the next step we derive the evolution equations for the amplitudes by projecting the full equation (2) onto the $\phi_{L}$ and $\phi_{R}$ modes using the ansatz (12)
$i\,\partial_{t}\big(c_{L}\phi_{L}+c_{R}\phi_{R}\big)=H\big(c_{L}\phi_{L}+c_{R}\phi_{R}\big).$ (14)
Multiplying on the left by $\phi_{L}(x)$ and integrating over $x$ we obtain an equation for $c_{L}$ (using orthonormality). The equation for $c_{R}$ is obtained in a similar way. The result is a pair of coupled linear ordinary differential equations (ODE)
$i\,\dot{c}_{L}$ $=$ $\langle\phi_{L}|H|\phi_{L}\rangle c_{L}+\langle\phi_{L}|H|\phi_{R}\rangle c_{R},$ (15)
$i\,\dot{c}_{R}$ $=$ $\langle\phi_{R}|H|\phi_{L}\rangle c_{L}+\langle\phi_{R}|H|\phi_{R}\rangle c_{R}.$ (16)
The matrix elements are calculated as follows
$\langle\phi_{L}|H|\phi_{L}\rangle=\langle\phi_{R}|H|\phi_{R}\rangle=\frac{\mu_{0}+\mu_{1}}{2},\quad\langle\phi_{L}|H|\phi_{R}\rangle=\langle\phi_{R}|H|\phi_{L}\rangle=\frac{\mu_{0}-\mu_{1}}{2}\;=\;\frac{\Delta}{2}.$ (17)
Thus, the matrix representation of $H$ in the R/L basis is given by
$H^{(LR)}=\left(\begin{array}[]{cc}(\mu_{0}+\mu_{1})/2&\Delta/2\\ \Delta/2&(\mu_{0}+\mu_{1})/2\\ \end{array}\right).$
The diagonal elements $(\mu_{0}+\mu_{1})/2$ only contribute to a global phase factor and can be ignored since they cancel in the calculation of probabilities and expectation values of observable quantities. Now, defining the tunneling amplitude $J=\Delta/2>0$, we can rewrite the Eqs. (15)-(16) as follows
$i\frac{d}{dt}\left(\begin{array}[]{c}c_{L}\\ c_{R}\end{array}\right)=\left(\begin{array}[]{cc}0&J\\ J&0\end{array}\right)\left(\begin{array}[]{c}c_{L}\\ c_{R}\end{array}\right).$
This yields the pair of coupled first-order ODEs
$i\,\dot{c}_{L}(t)=J\,c_{R}(t),\qquad i\,\dot{c}_{R}(t)=J\,c_{L}(t),$ (18)
where dots denote time derivatives. The final second-order ODE for $c_{L}$ is
$\ddot{c}_{L}(t)+J^{2}\,c_{L}(t)=0.$ (19)
This is a simple harmonic oscillator ODE with a frequency $\Omega=J$. A similar equation governs the other component $c_{R}(t)$. The general solution of Eq. (19) is
$c_{L}(t)=A\cos\!\left(\Omega t\right)+B\sin\!\left(\Omega t\right),$
with complex constants $A,B$ to be fixed by initial conditions. The probabilities to find the atom in the left and right wells of the DW potential are $P_{L}(t)=|c_{L}(t)|^{2}$ and $P_{R}(t)=|c_{R}(t)|^{2}$, respectively. Obviously, the normalization condition $P_{L}(t)+P_{R}(t)=1$ holds.
In analogy with Josephson oscillations we define the population imbalance $z(t)=P_{L}(t)-P_{R}(t)=1-2P_{R}(t)$. Using the special initial condition $c_{L}^{(0)}=1,c_{R}^{(0)}=0$ we obtain
$z(t)=\cos(\Omega t),\qquad\mbox{with}\qquad\Omega=\Delta/2.$ (20)
Figure 3 illustrates the real-time dynamics of the coupled system (1)–(2) for the parameter set $g=10$, $q=1$, $\gamma=3$, and $N=10$. Panel (a) shows the stationary double-well potential $|\psi(x,t)|^{2}$ created by the dipolar two-soliton molecule. As expected, the molecular density remains practically unchanged during the evolution, confirming that the impurity atom has a negligible back action on the condensate. Panel (b) presents the spatiotemporal evolution of the density $|\phi(x,t)|^{2}$, where the impurity periodically tunnels between the left and right wells of the self-induced potential, maintaining its localization and phase coherence over long time scales. Panel (c) displays the corresponding population imbalance
$z(t)=\int_{x<0}|\phi(x,t)|^{2}\,dx-\int_{x>0}|\phi(x,t)|^{2}\,dx,$ (21)
which demonstrates clear Josephson-like oscillations with a well-defined period. The measured oscillation period from direct numerical integration, $T_{\mathrm{meas}}=223.506$, agrees remarkably well with the theoretical prediction obtained from the two-mode model, $T_{\mathrm{pred}}=223.125$, in accordance with Eq. (20). This excellent agreement validates the analytical approximation and confirms that the impurity atom behaves as a coherent two-level system exhibiting Rabi-type oscillations between the two potential wells.
Figure 3: (a) The DW potential $\sim|\psi(x,t)|^{2}$ created by the dipolar two-soliton molecule remains stationary as time advances. (b) The density plot for $|\phi(x,t)|^{2}$ shows regular oscillations of the probability of finding the impurity atom on the left/right well. (c) The corresponding population imbalance $z(t)=\int_{x<0}|\phi(x,t)|^{2}dx-\int_{x>0}|\phi(x,t)|^{2}dx$ demonstrates coherent Josephson-type oscillations. The numerically measured oscillation period $T_{\mathrm{meas}}=223.506$ agrees closely with the theoretical prediction $T_{\mathrm{pred}}=223.125$ obtained from the two-mode model [Eq. (20)]. The parameters used are $g=10$, $q=1$, $w=5$, $\gamma=3$, and $N=10$.
V Tuning the qubit’s frequency
The ability to change the qubit’s frequency is crucial for its precise control and manipulation. Since the qubit’s frequency is determined by the energy difference $\Delta$ between its two lowest quantum states $\{\phi_{0},\phi_{1}\}$, the goal can be achieved by manipulating the underlying DW potential $V(x)\sim-|\psi(x)|^{2}$ in which the impurity atom is trapped. The DW potential can be modified by tuning the strength of atomic interactions in BEC, either short-range contact interactions $q$, or long-range dipolar interactions $g$. The variation of these parameters will result in corresponding variations of the distance $\sim a_{0}$ between the two wells of the DW potential, thus the desired energy splitting $\Delta=\mu_{1}-\mu_{0}$. When the two potential wells are sufficiently distant, the energy splitting disappears as the impurity atom’s tunneling probability approaches zero. In Fig. 4 we show the qubit’s frequency as a function of the strength of contact and dipolar atomic interactions. The curves demonstrate the tendency for the energy splitting $J=(\mu_{1}-\mu_{0})/2$ (proportional to the qubit’s oscillation frequency) to increase, as the two wells of the DW potential approach each other (width parameter $a_{0}$ decrease). Evidently, the effect of contact interactions is almost linear, while the long range dipolar interactions show hyperbolic behavior.
$\qquad\qquad(a)$                       $(b)$
Refer to caption  Refer to caption
Figure 4: The frequency of coherent oscillations of the impurity atom between the two wells of the double-well (DW) potential as a function of the interaction strengths in the dipolar condensate. (a) Dependence of the oscillation frequency on the contact interaction parameter $q$ for fixed $g=10$, $w=5$, $\gamma=3$, and $N=10$. (b) Dependence on the dipolar (nonlocal) interaction strength $g$ for fixed $q=1$, $w=5$, $\gamma=3$, and $N=5$. Blue circles represent the numerically obtained oscillation frequencies, while the yellow dashed curves correspond to the analytical predictions based on the variationally obtained width parameter $a_{0}$ from Eq. (8).
The dynamics of a qubit are best represented using the Bloch sphere. The qubit, initialized in the $\phi_{0}$ state, precesses around the XY-plane of the Bloch sphere. To create superposition states like $(\phi_{0}+\phi_{1})/\sqrt{2}$, it is necessary first to initialize the qubit in the state $\phi_{0}$, and then apply a Hadamard gate.
Figure 5 shows the Bloch sphere, which is generated directly from numerical simulations. We first construct the DW potential due to the dipolar two-soliton molecule, using the imaginary time propagation method applied to the governing GPE (1). Then, using the obtained DW potential, we solve the eigenvalue problem for Eq. (2) to find the quantum states of the impurity atom $\{\phi_{0}(x),\phi_{1}(x)\}$. Next, inserting these as initial conditions, we solve the coupled system (1)-(2) in real time. At each recorded time we project the impurity state onto the two basis modes, thus finding a pair of complex amplitudes. After re-normalizing, to ensure an exact two-level system, we converted those amplitudes onto the Bloch coordinates $r_{x}=2\,{\rm Re}(c_{0}c_{1}^{\ast}),\quad r_{y}=2\,{\rm Im}(c_{0}c_{1}^{\ast}),\quad r_{z}=|c_{0}|^{2}-|c_{1}|^{2}$. That gives one point on the sphere per time sample, forming the trajectory. Then, we plotted the light sphere, drew the trajectory, and marked the start/end points.
Refer to caption  Refer to caption
Figure 5: Left panel: The Bloch sphere represents a qubit state. The north pole corresponds to the ground state $\phi_{0}$ and the south pole corresponds to the first excited state $\phi_{1}$. An arbitrary superposition of these basis states can be obtained by parametrization $\phi=\cos(\theta/2)|\phi_{0}\rangle+e^{i\varphi}\sin(\theta/2)|\phi_{1}\rangle$. Right panel: The Bloch components $r_{x}\simeq\cos(\Delta t)$, $r_{y}\simeq\sin(\Delta t)$, $r_{z}\simeq 0$ show circular equatorial motion (upper). The purity of the two-level system calculated as ${\rm Tr}(\rho^{2})$, with $\rho$ being the density matrix, remains very close to unity indicating negligible de-coherence (lower). The parameters used are $g=10$, $q=1$, $w=5$, $\gamma=3$, and $N=10$.
VI Conclusions
We demonstrated that a single impurity atom embedded in a dipolar two-soliton molecule behaves as a controllable two-level system undergoing coherent Josephson type oscillations which are fundamental for the realization of qubits. We employed a variational approach to generate a two-soliton molecular potential, allowing practical control over qubit characteristics, including barrier height and tunneling frequency. The system is shown to be a good candidate for a qubit with two-level system’s purity ${\rm Tr}(\rho^{2})>0.9999$, as shown in Fig. 5. Future research should explore entanglement effects and potential decoherence mechanisms. The experimental realization of such a qubit using chromium, dysprosium or erbium dipolar BEC systems appears feasible within current ultracold-atom technologies.
Acknowledgements
We gratefully acknowledge the support provided by King Fahd University of Petroleum and Minerals (KFUPM) and the Research, Development and Innovation Authority (RDIA) for funding this work through project No. 22715-KFUPM-2023-KFUPM-B-4-1-EI.